OpenStudy (anonymous):

After a sailboat had been on the water for 5 h, a change in the wind direction reduced the average speed of the boat by 6 mph. The entire distance sailed was 94 mi. The total time spent sailing was 8 h. How far did the sailboat travel in the first 5 h?

OpenStudy (anonymous):

Let the distance traveled in first 5 hrs be X m. Then the average velocity in those 5 h = X/5 mph. Now, due to wind the velocity recudes by 6 mph i.e. the velocity now becomes : X/5 - 6 mph The boat sails with this new velocity for 8h-5h (Total hours - First 5h) = 3h so the distance traveled in this 3h duration is : (X/5 - 6) * 3 The total distance is 94 m i.e. X + (X/5 - 6)*3 = 94 : The total distance traveled is the distance traveled in first five hrs i.e. X (what we assumed) plus the distance traveled in next 3hrs i.e. (X/5 - 6)*3 (what we calculated) and that is equal to 94 m (given in the question. solve for X and you get the answer

OpenStudy (anonymous):

I still don't get it, what do I need to do.

OpenStudy (anonymous):

The idea is to assume any one quantity which is closest to the question asked then build on the equations based on the data given in the questions using the standard relation based on the newtons laws of motion.....solve the equations to get the unknown and finally get the question asked...